suppose it takes 3 days for space vehicle to travel from earth to the moon. about how long would it take the same vehicle traveling at the same speed to reach venus?
closest distance from earth to:
moon 240,000 mi
venus 26,000,000mi

Answer by stanbon(75887)

suppose it takes 3 days for space vehicle to travel from earth to the moon. about how long would it take the same vehicle traveling at the same speed to reach venus?
closest distance from earth to:
moon 240,000 mi
venus 26,000,000mi
------------------------------
Let the time neede to get to venus be "x".
Equation:
x/26,000,000 = 3/240,000
x = (3/240,000)*26,000,000
x = 325 days
===================
Cheers,
Stan H.